OpenStudy (anonymous):

suppose the limit of f(x)=f(4) as x approaches 4 then.. a) f is differentiable at x=4 b) f is continuous at x=4 c) f is continuous and differentiable at x=4 d) f is differentiable but not necessarily continuous at x=4

OpenStudy (anonymous):

c) f is continuous and differentiable at x=4

OpenStudy (turingtest):

a function is continuous at a point \(x=a\) iff\[\lim_{x\to a}f(x)=f(a)\]a function is differentiable at \(x=a\) iff\[\lim_{b\to a}\frac{f(a)-f(b)}{a-b}=f'(a)\]

OpenStudy (turingtest):

you are not given any information about the second criteria, hence...?

OpenStudy (anonymous):

it is continuous?

OpenStudy (turingtest):

yes